# Amazon Web Service

### v1

@@ 4 From Arithmetic to Relational Reasoning
Relational reasoning provides the perfect test case. Like arithmetic, it demands automatic binding from natural language to symbolic structure, followed by systematic execution of transformations. Just as “9.11” must be bound as a numerical quantity rather than a historical date, “Alice” must be bound to appropriate relational roles before applying logical transformations like $\forall x,y$ [Parent$(x,y)\rightarrow$ Child$(y,x)$].
@@
Computational Execution Execute bound operations: $9.9>9.11$ Execute inference rules: Parent$(X,Y)\rightarrow$ Child$(Y,X)$
Consistency Maintain across multi-step procedures: $a>b\land b>c\rightarrow a>c$ Maintain variable bindings across multi-step inference: Parent$(A,B)\land$ Parent$(B,C)\rightarrow$ Grandparent$(A,C)$
@@ 4.1 The Reversal Curse and Alice Problem
Consider the computational requirements for the reversal curse: given “Tom Cruise’s mother is Mary Lee Pfeiffer” and asked “Who is Mary Lee Pfeiffer’s son?”, models must automatically infer that the tokens represent PERSON entities, bind the implicit MOTHER_OF relationship from surface syntax, recognize the query seeks the inverse relationship, apply the symmetric transformation Mother$(x,y)\rightarrow$ Son$(y,x)$ in reverse, bind variables $x=$Mary Lee, $y=$Tom, and conclude Tom is Mary Lee’s son. Similarly, the Alice problem requires binding family structure from “has 4 sisters and 1 brother,” then rebinding Alice from SISTER role to SIBLING role to execute the perspective-shift computation.
@@ 4.2 Rule Applications, Logical Inconsistencies and Operator Failures
(Vashishtha et al., 2025) demonstrates success on causal reasoning tasks that test rule applications without variable binding. Pattern matching works when models are trained from scratch on clean synthetic data, with a 67M-parameter transformer outperforming billion-parameter LLMs on causal axioms. This confirms that direct rule application falls within transformers’ capabilities when training conditions isolate logical operations from the contextual contamination of multi-domain pretraining.
@@
The comprehensive study by Li et al. (2024) provides decisive evidence of LLMs’ architectural limitations in relational reasoning. They systematically evaluated state-of-the-art LLMs on inductive logic programming (ILP) tasks across both natural language and truth-value matrix formats, comparing performance against Differentiable Logic Machines (DLM) (Zimmer et al., 2023)—specialized neural program induction models designed explicitly for logical reasoning.
@@ 4.3 The Pattern Across Relational Reasoning
Our analysis reveals a convergent pattern across complexity levels: LLMs fail systematically regardless of how relational reasoning tasks are presented. Problems that appear simple (Reversal Curse, Alice problem) actually require sophisticated computational transformations that LLMs cannot execute. Problems that appear complex but provide explicit logical structure (truth-value matrices, clear logical relationships) should be easier to solve, yet LLMs still fail systematically. Table 2 summarizes the magnitude of these systematic failures, revealing consistently poor performance despite the apparent simplicity of many tasks.
@@
The performance gap persists even when LLMs are given identical inputs to specialized models: despite being 100,000 times larger than DLM models, LLMs performed dramatically worse across all logical reasoning tasks (Li et al., 2024), demonstrating that architectural constraints, not data availability, prevent reliable logical reasoning. This provides definitive evidence that the failures reflect fundamental limitations in how transformers process symbolic relationships, confirming our architectural analysis from Section 3.
@@ 5.1 Self-Scaffolding: Leveraging Comprehension for Step-by-Step Execution
“To multiply $742\times 89$: First, multiply $742\times 9=6678$. Then multiply $742\times 80=59360$. Finally, add $6678+59360=66038$.”
@@ 5.2 Tool Delegation: Bypassing Model Execution Entirely
The comment on the Illusion of Thinking paper (Opus and Lawsen, 2025) provides a particularly striking example of tool delegation’s effectiveness. When Tower of Hanoi problems were reformulated as: “Solve Tower of Hanoi with 15 disks. Output a Lua function that prints the solution when called,” models achieved high accuracy across tested systems (Claude-3.7-Sonnet, Claude Opus 4, OpenAI o3, Google Gemini 2.5), completing in under 5,000 tokens. This dramatic improvement occurred because models generated algorithmic code rather than attempting to enumerate moves themselves. Similarly, Cheng et al. (Cheng et al., 2024) demonstrate that separating symbolic code generation from symbolic code execution improves deductive reasoning.
@@ 5.3 Hybrid Architectures: Specialized Modules for Symbolic Operations
Current Mixture-of-Experts (MoE) architectures optimize primarily for computational efficiency—routing tokens to different parameter sets to increase model capacity—rather than functional specialization based on cognitive capabilities. A neuroscience-inspired approach could allocate specific experts to handle symbolic operations, relational reasoning, pattern completion, and other specialized functions. Just as the human brain has dedicated regions for executive reasoning (prefrontal cortex) and memory (temporal lobe), functionally partitioned MoE could develop experts that excel at exactly the symbolic binding and manipulation operations that current general-purpose feed-forward networks struggle to implement.
@@ 5.4 Coordination and The Metacognitive Bottleneck
1:  $P\leftarrow$ input problem
2:  $d\leftarrow$ assess_difficulty$(P)$requires introspection}
@@
4:     return pattern_complete$(P)$
@@
6:     $P^{\prime}\leftarrow$ decompose_or_delegate$(P)$tool call / planner}
7:     return solve$(P^{\prime})$
@@
To test the importance of metacognition, we designed a tightly controlled experiment using $n$-digit × $n$-digit multiplication tasks with place-value decomposition and simulated column-wise addition, and tested Claude Sonnet 4, GPT-4o, and Gemini 2.5 Flash across three conditions:
@@
Table 3: Performance on multiplication problems with golden decomposition prompts
@@ 6 Pattern Completion: Strengths and Limitations
The results reveal a fundamental performance cliff: while OpenAI’s o3 achieved 87.5% on ARC-AGI-1 through sophisticated pattern completion, the same architecture scored $<3\%$ on ARC-AGI-2 (Chollet et al., 2025; ARC Prize Foundation, 2025). Performance plummets once pattern completion loses its near-neighbor anchors from training, despite identical instruction formats. This cliff occurs precisely where tasks transition from pattern recognition to genuine rule discovery and systematic reasoning—exposing the architectural limitations we have identified throughout this paper.
@@
However, discovery of novel algorithms such as solving the Tower of Hanoi problem from scratch requires a much stronger set of capabilities. The optimal algorithm that works for any number of disks $n$ requires recognizing that moving $n$ disks involves first moving $n-1$ disks, then the largest disk, then the $n-1$ disks again. This recursive insight, itself a meta-pattern, is even more challenging than what ARC-AGI-2 tests. Recent evaluation shows that state-of-the-art reasoning models like Claude 3.7 Sonnet and OpenAI’s o3-mini exhibit complete performance collapse beyond 8-9 disks, despite having sufficient token budget to continue (Shojaee et al., 2025).
@@ 7 Conclusion
This diagnosis clarifies why interpretability tools often uncover expressive artifacts rather than reliable algorithms, and why model self-explanations may diverge from the actual pathways used in computation. It also explains how training order can influence which internal patterns are embedded and where—even when outward behavior remains stable—causing interpretability efforts to surface path-dependent artifacts rather than consistent mechanisms. These insights reframe current conversations about emergence and scaling: more data or parameters cannot resolve bottlenecks rooted in architectural design.
@@ A.2 Impossibility of Exact Multiplication
Appendix B Metacognitive Assessment Experiments
B.1 Experimental Design

### v2

@@ 4 From Arithmetic to Relational Reasoning
[…]
@@
[…]
Consistency Maintain across multi-step procedures: $a>b\land b>c\rightarrow a>c$ Maintain variable bindings across multi-step inference: $(A,B)\land Parent$$(B,C)\rightarrow$ Grandparent(A,C)(A,C)
@@ 4.1 The Reversal Curse and Alice Problem
[…]
@@ 4.2 Rule Applications, Logical Inconsistencies and Operator Failures
Vashishtha et al. (2025) demonstrates success on causal reasoning tasks that test rule applications without variable binding. Pattern matching works when models are trained from scratch on clean synthetic data, with a 67M-parameter transformer outperforming billion-parameter LLMs on causal axioms. This confirms that direct rule application falls within transformers’ capabilities when training conditions isolate logical operations from the contextual contamination of multi-domain pretraining.
@@
The comprehensive study by Li et al. (2024a) provides decisive evidence of LLMs’ architectural limitations in relational reasoning. They systematically evaluated state-of-the-art LLMs on inductive logic programming (ILP) tasks across both natural language and truth-value matrix formats, comparing performance against Differentiable Logic Machines (DLM) (Zimmer et al., 2023)—specialized neural program induction models designed explicitly for logical reasoning.
@@ 4.3 The Pattern Across Relational Reasoning
Our analysis reveals a convergent pattern across complexity levels: LLMs fail systematically regardless of how relational reasoning tasks are presented. Problems that appear simple (Reversal Curse, Alice problem) actually require sophisticated computational transformations that LLMs cannot execute. Problems that appear complex but provide explicit logical structure (truth-value matrices, clear logical relationships) should be easier to solve, yet LLMs still fail systematically. Table 2 summarizes the magnitude of these systematic failures (data from (Berglund et al., 2023; Nezhurina et al., 2024; Li et al., 2024a)), revealing consistently poor performance despite the apparent simplicity of many tasks.
@@
The performance gap persists even when LLMs are given identical inputs to specialized models: despite being 100,000 times larger than DLM models, LLMs performed dramatically worse across all logical reasoning tasks (Li et al., 2024a), demonstrating that architectural constraints, not data availability, prevent reliable logical reasoning. This provides definitive evidence that the failures reflect fundamental limitations in how transformers process symbolic relationships, confirming our architectural analysis from Section 3.
@@ 5.1 Self-Scaffolding: Leveraging Comprehension for Step-by-Step Execution
“To multiply $742\times 89$: First, multiply $742\times 9=6678$. Then multiply $742\times 80=59360$. Finally, add 6678+59360=660386678+59360=66038.”
@@ 5.2 Tool Delegation: Bypassing Model Execution Entirely
The necessity of such delegation becomes clear when we consider the structured data operations analyzed in Section 3.4. Operations such as finding maximum values or computing averages require iterative state tracking that LLMs cannot perform reliably, yet these same models excel at recognizing when to invoke SQL or pandas functions that implement these algorithms correctly. This pattern extends beyond data operations to other domains requiring algorithmic execution.
The comment on the Illusion of Thinking paper (Opus and Lawsen, 2025) provides a particularly striking example of tool delegation’s effectiveness. When Tower of Hanoi problems were reformulated as: “Solve Tower of Hanoi with 15 disks. Output a Lua function that prints the solution when called,” models achieved high accuracy across tested systems (Claude-3.7-Sonnet, Claude Opus 4, OpenAI o3, Google Gemini 2.5), completing in under 5,000 tokens. This dramatic improvement occurred because models generated algorithmic code rather than attempting to enumerate moves themselves. Similarly, Cheng et al. (2024) demonstrate that separating symbolic code generation from symbolic code execution improves deductive reasoning.
@@ 5.3 Hybrid Architectures: Specialized Modules for Symbolic Operations
Current Mixture-of-Experts (MoE) architectures (Shazeer et al., 2017) optimize primarily for computational efficiency—routing tokens to different parameter sets to increase model capacity—rather than functional specialization based on cognitive capabilities. A neuroscience-inspired approach could allocate specific experts to handle symbolic operations, relational reasoning, pattern completion, and other specialized functions. Just as the human brain has dedicated regions for executive reasoning (prefrontal cortex) and memory (temporal lobe) (Kandel et al., 2021), functionally partitioned MoE could develop experts that excel at exactly the symbolic binding and manipulation operations that current general-purpose feed-forward networks struggle to implement.
@@ 5.4 Coordination and The Metacognitive Bottleneck
[…]
@@
4:  return pattern_complete(P)(P)
@@
6:  $P^{\prime}\leftarrow$ decompose_or_delegate(P)(P) {tool call / planner}
[…]
@@
To test the importance of metacognition, we designed a tightly controlled experiment using $n$-digit × $n$-digit multiplication tasks with place-value decomposition and simulated column-wise addition. We tested Claude Sonnet 4, GPT-4o, and Gemini 2.5 Flash across three conditions, with results summarized in Table 3:
@@
Table 3: Performance on multiplication problems with golden decomposition prompts. Sum errors refers to failures in final addition despite correct intermediate steps.
@@ 6 Pattern Completion: Strengths and Limitations
The results reveal a fundamental performance cliff: while OpenAI’s o3 achieved 87.5% on ARC-AGI-1 through sophisticated pattern completion, the same architecture scored $3\%$ on ARC-AGI-2 (Chollet et al., 2025; ARC Prize Foundation, 2025). Performance plummets once pattern completion loses its near-neighbor anchors from training, despite identical instruction formats. This cliff occurs precisely where tasks transition from pattern recognition to genuine rule discovery and systematic reasoning—exposing the architectural limitations we have identified throughout this paper.
@@
However, discovery of novel algorithms such as solving the Tower of Hanoi problem from scratch requires a much stronger set of capabilities. The optimal algorithm that works for any number of disks $n$ requires recognizing that moving $n$ disks involves first moving n1n-1 disks, then the largest disk, then the n1n-1 disks again. This recursive insight, itself a meta-pattern, is even more challenging than what ARC-AGI-2 tests. Recent evaluation shows that state-of-the-art reasoning models like Claude 3.7 Sonnet and OpenAI’s o3-mini exhibit complete performance collapse beyond 8-9 disks, despite having sufficient token budget to continue (Shojaee et al., 2025).
@@
7 Implications for Mechanistic Interpretability
Having established three architectural constraints that create the computational split-brain syndrome, we now examine how our framework relates to findings from mechanistic interpretability research. While these sophisticated tools have revealed important insights about model internals, we argue that the architectural constraints we identify suggest an alternative interpretation of what these tools discover: sophisticated pattern coordination rather than algorithmic implementation. This perspective does not diminish the value of interpretability research but rather reframes what we might be learning from it.
7.1 From Activation Patching to Sparse Autoencoders
Early mechanistic interpretability relied heavily on activation patching, intervening on specific neurons or attention heads to understand their causal role in model behavior. This approach revealed that arithmetic in LLMs rely not on clean algorithmic circuits but on what Nikankin et al. (2024) characterize as a “bag of heuristics”—distributed, overlapping pattern detectors with no coherent computational structure.
However, activation patching faces several limitations that motivated the development of more sophisticated methods. The polysemantic nature of individual neurons makes it difficult to isolate specific computational functions through patching alone—single neurons often represent multiple concepts simultaneously. Interventions that work on training-like inputs often fail on novel examples, revealing poor out-of-distribution generalization. Furthermore, our analysis suggests that the specific neurons and attention heads involved in any computation are highly path-dependent, varying significantly across training runs even when final behavior converges, making findings from activation patching potentially idiosyncratic to specific model instances.
Sparse Autoencoders (SAEs) emerged as a proposed solution to the polysemanticity problem (Bricken et al., 2023; Templeton et al., 2024). SAEs are unsupervised learning models that decompose neural network activations into sparse, interpretable features by learning an overcomplete dictionary of feature directions. By expanding activations into a higher-dimensional space with sparsity constraints, SAEs aim to identify monosemantic units—features that capture single, interpretable concepts rather than the mixed representations found in individual neurons. These features appear more stable across models, with Wang et al. (2025) demonstrating that similar features emerge across different open-source architectures trained on comparable data.
Recent methods extend interpretability analysis from local interventions to whole-network computation traces. Linear Computation Graphs (He et al., 2024) build on SAE features to trace complete computational pathways across layers, while Attribution Graphs (Lindsey et al., 2025a) trace information flow through raw model activations without requiring feature decomposition. Despite their different approaches—one working through SAE features, the other through raw activations—both methods reveal similar patterns of distributed, redundant computation rather than clean algorithmic circuits.
7.2 Alternative Interpretations of Interpretability Findings
The interpretability findings discussed above are typically interpreted as evidence of computational mechanisms within LLMs. Our architectural framework, however, suggests an alternative lens through which to view these same empirical results—one that emphasizes pattern coordination over algorithmic implementation. This interpretation, while consistent with the data, differs from how the original authors might characterize their findings.
Consider the Indirect Object Identification (IOI) circuit identified through Linear Computation Graphs (He et al., 2024). The circuit traces how models process sentences like “When Mary and John went to the store, John gave Mary a bottle,” identifying features that track entities and their relationships. Yet, our framework suggests these features recognize statistical patterns—names appearing early in sentences often receive objects mentioned later—rather than implementing genuine reference resolution through variable binding. The original authors might interpret these circuits as genuine computational mechanisms; our framework offers an alternative interpretation where these represent sophisticated statistical pattern recognition.
The cross-architecture universality of features (Wang et al., 2025) initially seems to suggest discovery of fundamental computational units. However, our framework suggests this universality could reflect convergent solutions to pattern matching. Whether these represent algorithmic primitives or statistical patterns remains an open empirical question. Models trained on similar corpora develop similar pattern-recognition strategies because they encounter similar statistical regularities.
This interpretation aligns with the “bag of heuristics” finding (Nikankin et al., 2024). While their analysis examined neuron-level patterns rather than SAE features, our framework predicts—though this remains to be empirically tested—that similar limitations would apply to SAE features. The heuristics are not implementation accidents that better interpretability tools could resolve into clean algorithms. They reflect the fundamental strategy available to pattern-completion architectures: coordinate multiple pattern detectors to approximate computational outcomes without implementing computational procedures. SAE features may be more stable across training runs than individual neurons—they capture statistical regularities that any model trained on similar data must learn. But these stable regularities remain patterns rather than algorithms.
The apparent success of mechanistic interpretability may itself exemplify the computational split-brain syndrome. We excel at comprehending what features correlate with—recognizing that certain features activate on arithmetic, logical relations, or syntactic structures. Yet, we may be mistaking this comprehension for understanding of computational competence. A feature that activates on multiplication and whose ablation disrupts multiplication could be recognizing multiplication patterns rather than implementing multiplication. The very interpretability of these features—that they correspond to human-recognizable concepts—might bias us toward assuming they implement rather than recognize these concepts.
7.3 Critical Questions for Interpretability Research
Our framework suggests two methodological directions that could help distinguish between pattern-matching and algorithmic hypotheses in interpretability research:
Systematic Testing at Scale.
Current interpretability work often demonstrates features on expected patterns without systematic out-of-distribution testing. Following the methodology of Nikankin et al. (2024) and GSM-Symbolic, interpretability claims should be validated across:
  • Surface form variations (“3 + 5” vs. “three plus five” vs. “add 3 and 5”)
  • Systematic perturbations (changing numbers, adding irrelevant clauses, reordering)
  • Distribution shifts that preserve computational requirements
If SAE features truly implement computational operations rather than pattern matching, they should maintain their function across these variations. Current evidence suggests they do not—features that appear to track arithmetic operations fail when numbers appear in unexpected formats, revealing pattern recognition rather than algorithmic implementation.
Addressing Confirmation Bias.
Interpretability research currently suffers from a methodological limitation: features are typically interpreted after observing their activation patterns on known examples. This post-hoc analysis invites confirmation bias—we see what we expect to see. A more rigorous approach would:
  • Interpret features based solely on their structure before examining behavior
  • Make explicit predictions about feature responses to novel inputs
  • Report both successful and failed predictions transparently
  • Test features on adversarial examples designed to separate correlation from causation
These methodological improvements would help distinguish between two fundamentally different types of findings:
  1. 1.
    Statistical Mechanism Discovery: Understanding how models coordinate pattern matching to produce outputs—what current methods achieve effectively
  2. 2.
    Algorithmic Mechanism Discovery: Identifying computational procedures that generalize beyond training distributions—what may be architecturally impossible in current transformers
We emphasize that current interpretability research has made substantial contributions to understanding how LLMs process information. Our critique is not of the methods or findings themselves, but rather suggests a reframing of what these findings might represent given the architectural constraints we identify. By acknowledging that we study sophisticated pattern coordination rather than algorithmic implementation, interpretability research can provide more accurate insights into both the capabilities and fundamental limitations of current language models. This reframing may ultimately prove more valuable than maintaining the illusion that we are reverse-engineering computational mechanisms that may not exist.
8 Reflections on the Research Journey
This work emerged from a multi-year investigation that began with a seemingly simple question: how do LLMs perform grade-level math problems that, while conceptually simple, require multi-step algorithmic execution? The search for generalizable arithmetic circuitry led to a deeper puzzle—why should grade-school mathematics ever present an “out-of-distribution” problem for models trained on vast corpora? In parallel, the Reversal Curse and Alice-in-Wonderland problems revealed that automatic binding to concepts and relations—something humans take for granted—simply does not occur in LLMs. The viral “9.11 vs 9.9” comparison, initially dismissed as an amusing anecdote, crystallized as a fundamental input representation problem: tokens carry contextual contamination that prevents clean symbolic binding. Gradually, we recognized these weren’t isolated curiosities but systematic failures spanning arithmetic, logic, and relational reasoning—all stemming from the same architectural constraints. This journey from searching for arithmetic circuits to discovering their impossibility ultimately revealed the computational split-brain syndrome as a unifying explanation for why LLMs can explain what they cannot execute.
@@ 9 Conclusion
This diagnosis clarifies why interpretability tools often uncover expressive artifacts rather than reliable algorithms, and why model self-explanations may diverge from the actual pathways used in computation. Our finding that instruction and execution pathways are geometrically separated raises fundamental concerns: models’ explanations of their reasoning may not reflect their actual computational processes, challenging current approaches to explainability. Training order influences which internal patterns are embedded and where—even when outward behavior remains stable—causing interpretability efforts to surface path-dependent artifacts rather than consistent mechanisms.
These insights reframe current conversations about emergence and scaling. Our findings align with critiques by Bender et al. (2021) and others who argue that the "bigger-is-better" paradigm has fundamental flaws. By demonstrating that LLM limitations are architectural rather than scale-dependent, we provide technical evidence that more data or parameters cannot resolve bottlenecks rooted in architectural design. This suggests redirecting resources from pure scaling toward architectural innovations could reduce computational and environmental costs while improving reliability.
@@
Acknowledgments
This work benefited greatly from insightful discussions and valuable feedback from colleagues. Special thanks to Yasser Shaaban, David Paul Wipf, Tong He, and many students for reviewing early drafts and offering constructive comments. Tong He is also gratefully acknowledged for assistance in refining the experiments. We thank the TMLR reviewers for their thoughtful feedback, which substantially improved the clarity and rigor of this work.
@@ A.2 Impossibility of Exact Multiplication
Appendix B Appendix: Role Specialization Under Next-Token Prediction
We provide a detailed proof sketch for the role specialization theorem presented in Section 3.2.
B.1 Setup and Notation
Consider a transformer with $L$ layers trained on next-token prediction. We denote:
  • $\theta=\{\theta_{E},\theta_{A},\theta_{F}\}$: The three parameter sets for embeddings, attention, and FFNs respectively
  • $\mathcal{D}$: Training distribution over token sequences
  • $\mathcal{L}(\theta)=\mathbb{E}_{(w_{1},...,w_{T})\sim\mathcal{D}}[-\log p_{\theta}(w_{t}|w_{<t})]$: The expected cross-entropy loss
  • $x^{(L)}\in\mathbb{R}^{d}$: Final layer representation
  • $E\in\mathbb{R}^{|V|\times d}$: Embedding/unembedding matrix
  • $p_{\theta}(w_{t}|w_{<t})=\text{softmax}(x^{(L)}\cdot E^{T})[w_{t}]$: Next-token probability
B.2 Gradient Flow Analysis
For arithmetic tasks like “43 × 78 = 3354”, the model must minimize:
$\mathcal{L}_{\text{local}}=-\log p_{\theta}(\text{`3354'}|\text{`43 \texttimes 78 ='})$
This requires $x^{(L)}\cdot e_{3354}>x^{(L)}\cdot e_{v}$ for all $v\neq 3354$ in the vocabulary.
Gradient w.r.t. Final Representation.
The gradient of the loss with respect to the final representation is:
$\frac{\partial\mathcal{L}}{\partial x^{(L)}}=E^{T}(p_{\theta}-y_{\text{onehot}})$
where $p_{\theta}$ is the predicted distribution and $y_{\text{onehot}}$ is the one-hot encoding of the target token.
Backpropagation Through Components.
Since $x^{(L)}=\text{embeddings}+\sum_{l=1}^{L}\Delta h^{(l)}_{\text{attn}}+\sum_{l=1}^{L}\Delta h^{(l)}_{\text{FFN}}$, the gradient flows to:
  1. 1.
    Embeddings ($\theta_{E}$):
    $\frac{\partial\mathcal{L}}{\partial\theta_{E}}=\frac{\partial\mathcal{L}}{\partial x^{(L)}}\cdot\frac{\partial x^{(L)}}{\partial\text{embeddings}}\cdot\frac{\partial\text{embeddings}}{\partial\theta_{E}}$
    But embeddings must serve multiple contexts. For token “43”, the embedding must minimize loss across:
    • Mathematical contexts: “43 + 57”, “43 × 2”
    • Date contexts: “43 BC”, “1943”
    • Version contexts: “version 43.0”
    The multi-objective optimization forces:
    $\theta_{E}^{*}(\text{``43''})=\arg\min_{e}\sum_{c\in\mathcal{C}(\text{``43''})}\mathcal{L}_{c}(e)$
    This averaging across contexts prevents the isometric properties needed for arithmetic (see Section 3.1).
  2. 2.
    Attention ($\theta_{A}$): At the “=” position with causal masking, attention computes:
    $\Delta h_{\text{attn}}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$
    where $Q$, $K$, $V$ are projections of $\{$“43”, “×”, “78”, “=”}\}.
    Key Constraint: The output is a weighted average:
    $\Delta h_{\text{attn}}\in\text{conv}\{v_{1},v_{2},v_{3},v_{4}\}$
    where $v_{i}$ are the value vectors. Since $e_{3354}\notin\text{conv}\{e_{43},e_{\times},e_{78},e_{=}\}$, attention cannot generate the required output direction. It can only create rich encodings of the operation context.
  3. 3.
    FFNs ($\theta_{F}$): FFNs must map attention’s encoding to the result. The required function is:
    $F:h_{\text{attn}}(\text{``multiply 43 by 78''})\mapsto\text{direction}(e_{3354})$
    By UAT, FFNs with sufficient width can approximate any continuous function on compact sets. Define:
    $K=\{h_{\text{attn}}(a\times b):(a,b)\in\text{training data}\}$
    For any $\epsilon>0$, there exist weights such that:
[…]
    Critical Issue: This approximation only holds on $K$. For novel arithmetic outside $K$, no approximation guarantee exists.
B.3 Why Memorization Is Inevitable
Architectural Impossibility.
FFNs implement $\text{FFN}(x)=W_{2}\cdot\text{ReLU}(W_{1}x+b_{1})+b_{2}$, which is piecewise linear. Multiplication $f(a,b)=a\times b$ is not piecewise linear:
$\frac{\partial^{2}f}{\partial a\partial b}=1\neq 0$
No weight configuration can make a piecewise linear function compute exact multiplication (see Theorem 2 in Appendix A).
Training Dynamics.
Given $N$ training examples $\{(a_{i}\times b_{i},c_{i})\}_{i=1}^{N}$, gradient descent optimizes:
$\theta_{F}^{*}=\arg\min_{\theta_{F}}\sum_{i=1}^{N}-\log p(\text{token}(c_{i})|h_{\text{attn}}(a_{i}\times b_{i}))$
Since exact computation is impossible, the optimization converges to memorizing the mapping:
$h_{\text{attn}}(a_{i}\times b_{i})\mapsto e_{c_{i}}$
This is pattern storage, not algorithmic learning. The UAT capacity that could theoretically approximate multiplication is instead allocated to memorizing these $N$ discrete mappings.
Out-of-Distribution Failure.
For novel multiplication $(a_{\text{new}},b_{\text{new}})\notin\text{training}$:
  • $h_{\text{attn}}(a_{\text{new}}\times b_{\text{new}})\notin K$
  • FFN has no memorized pattern
  • Output is interpolation between nearest memorized patterns
  • Result is essentially random from neighboring vocabulary tokens
B.4 Conclusion
The role specialization is inevitable under gradient descent:
  1. 1.
    Embeddings must average across diverse contexts, losing arithmetic structure
  2. 2.
    Attention can only compute weighted averages, unable to generate novel tokens
  3. 3.
    FFNs cannot implement multiplication algorithmically, so they memorize patterns
This specialization emerges not from insufficient capacity—FFNs have universal approximation capability—but from how the next-token prediction objective allocates that capacity. The result is pattern storage rather than algorithmic computation, explaining why LLMs can articulate algorithms perfectly while failing to execute them reliably. $\square$
Appendix C Metacognitive Assessment Experiments
C.1 Experimental Design